OpenStudy (anonymous):

What is the solution of the following system? 3x + 3y = 10 -9x - 9y = -30 please show steps

hero (hero):

-3(3x + 3y) = -3(10)

OpenStudy (***[isuru]***):

take the second one -9x - 9y = -30 divide each term by (-3) (-9x)/(-3) + (-9y)/(-3) = (-30)/(-3) where u will get 3x + 3y = 10 which is exactly same as the first one.. which means there r infinitely many solutions to this hope this will help ya!!

OpenStudy (anonymous):

ahh that makes sooo much more since thank you!!

OpenStudy (***[isuru]***):

u r welcome !

hero (hero):

@spikey419, I was trying to motivate you to think for yourself rather than spell out all the steps for you